Ok stuck on this one.
Jill has $3.50 in nickels and dimes. If she has 50 coins, how many of each type of coin does she have?

Found 2 solutions by scott8148, stanbon:
Answer by scott8148(6628)

let x=nickles so 50-x=dimes ... 5x+10(50-x)=350 ... 5x+500-10x=350 ... -5x=-150 ... x=30 ... 50-x=20

Answer by stanbon(75887)

Jill has $3.50 in nickels and dimes. If she has 50 coins, how many of each type of coin does she have?
----------
Let # of nickels be "x" ; Value of these is 5x cents.
# of dimes is "50-x" ; Value of these is 10*(50-x) = 500-10x cents
----------------
EQUATION:
value + value = 350 cents
5x + 500-10x = 350
-5x = -150
x=30 (number of nickels)
50-x = 20 (number of dimes)
============
Cheers,
Stan H.